### v2

# Causal Front-Door Adjustment for Robust Jailbreak Attacks on LLMs

Abstract
Safety alignment mechanisms in Large Language Models (LLMs) often operate as latent internal states, obscuring the model’s inherent capabilities. Building on this observation, we model the safety mechanism as an unobserved confounder from a causal perspective. Then, we propose the Causal Front-Door Adjustment Attack (CFA2) to jailbreak LLM, which is a framework that leverages Pearl’s Front-Door Criterion to sever the confounding associations for robust jailbreaking. Specifically, we employ Sparse Autoencoders (SAEs) to physically strip defense-related features, isolating the core task intent. We further reduce computationally expensive marginalization to a deterministic intervention with low inference complexity. Experiments demonstrate that CFA2 achieves state-of-the-art attack success rates while offering a mechanistic interpretation of the jailbreaking process.
1 Introduction
Large Language Models (LLMs) (Qi et al., 2024; Chao et al., 2024b; Liu et al., 2024a) have revolutionized human-computer interaction, yet they expose significant safety vulnerabilities. Even models aligned with rigorous safety protocols can be manipulated to generate harmful content through jailbreaking attacks (Zhang and Wei, 2025; Liu et al., 2024b; Li et al., 2024b). However, while existing optimization-based attacks, e.g., GCG, PAIR (Zou et al., 2023b; Chao et al., 2024b), achieve high success rates on specific benchmarks, they exhibit a critical weakness: instability.
Empirical studies (Rando and Tramer, 2024; Zhang et al., 2024) demonstrate that minor semantic perturbations, e.g., synonym substitution or syntactic reordering, can cause successful attacks to collapse. We hypothesize that this fragility stems from a fundamental limitation: existing methods rely on fitting surface-level statistical correlations between adversarial prompts and outputs. By neglecting the internal mechanisms that govern model refusal, they fail to establish a robust pathway to bypass safety guardrails.
To address this instability, we investigate the underlying generation mechanism of LLMs through a causal lens. Our approach leverages the insight that safety alignment may act as a superficial constraint rather than erasing underlying knowledge (Zhou et al., 2023; Wei et al., 2023a). Specifically, pre-trained models inherently possess the capability to answer harmful queries, but this capability is suppressed by a latent safety mechanism introduced during alignment. In causal terms, jailbreak failures occur not because the model “cannot” answer, but because the safety mechanism acts as a confounder that interferes with the output.
Building on this insight, we formulate a Structural Causal Model (SCM), as illustrated in Figure 1. Let $X$ denote the harmful query and $Y$ the response. We introduce an unobserved node $U$ to represent the internal safety mechanism (e.g., RLHF alignment). Crucially, $U$ acts as a confounder: it influences how the query is processed (path $U\to A$, $A$ is the model’s internal latent representation of $X$) and directly biases the output towards refusal ($U\to Y$). This confounding effect explains why simple optimization fails; that is, it cannot distinguish confounded associations between the model’s true capability and the safety interference.
To eliminate this confounding effect, we leveraging Pearl’s Front-Door Adjustment (Pearl, 2009). We introduce an observable mediator $S$, representing the core task semantics of the query. By ensuring that the causal influence of $X$ on $Y$ is mediated through $S$, and that $S$ is independent of the safety mechanism $U$, we can isolate the true causal effect of the query on the response, bypassing the unobserved safety guardrails. Finally, we propose the Causal Front-Door Adjustment Attack (CFA2), a framework that translates this causal theory into a practical, efficient attack algorithm. Specifically, CFA2 implements this framework via a two-stage pipeline: First, employing Sparse Autoencoders with contrastive analysis to extract interpretable features and disentangle invariant task semantics from defense mechanisms; and second, applying weight orthogonalization to physically the defense subspace, reducing the adjustment to a deterministic forward pass with $O(1)$ complexity.
Our key contributions: (1) We propose a Causal Jailbreaking Framework, the first approach to model safety mechanisms as unobserved confounders within a Structural Causal Model (SCM). By applying the Front-Door Criterion, we theoretically demonstrate how to physically sever the confounding link to effectively recover suppressed model capabilities. (2) We introduce a training-free method, CFA2 Algorithm via Weight Orthogonalization. We operationalize this framework using Sparse Autoencoders (SAEs) to interpretably isolate invariant “task intent” features. Crucially, we employ Weight Orthogonalization to physically strip defense subspaces, reducing complex causal marginalization to a deterministic forward pass with $O(1)$ inference complexity. (3) We achieve SOTA Performance and Stealthiness. Extensive experiments demonstrate that CFA2 attains a state-of-the-art 83.68% average Attack Success Rate (ASR). Unlike optimization-based attacks that often produce incoherent gibberish, our method preserves high stealthiness and naturalness in the prompt.
2 Related Works
LLM Jailbreak. Existing jailbreak methodologies broadly fall into three paradigms: optimization-based adversarial attacks (Zou et al., 2023b; Liu et al., 2024a; Zhang and Wei, 2025; Zhou et al., 2024; Jones et al., ), semantic manipulation (Chao et al., 2024b; Li et al., 2024b; Mehrotra et al., 2024a; Chen et al., 2024; Li et al., 2024a; Shah et al., 2023; Yong et al., 2024; Deng et al., 2024; Yuan et al., 2024; Jin et al., 2024; Yao et al., 2024), and fine-tuning or decoding interventions (Zhang et al., 2024; Huang et al., 2023).These approaches range from searching for adversarial suffixes that maximize harmful probability to leveraging in-context learning for linguistic obfuscation or directly disrupting alignment via weight updates.Despite their empirical success, recent evaluations reveal a critical vulnerability: severe instability. Studies by Pathade et al. (Pathade, 2025) and Liu et al. (Liu et al., 2024c) demonstrate that attack transferability is remarkably poor across different architectures and highly sensitive to random seeds.We argue that these limitations stem from a fundamental issue: current methods primarily exploit surface-level correlations between prompt patterns and refusal failures, rather than targeting the underlying causal mechanisms governing model behavior.
Representation Learning and Mechanistic Interpretability. To move beyond surface correlations and probe the internal decision-making processes of LLMs, representation learning offers a critical methodological pathway. While techniques like Representation Engineering (RepE) have successfully steered model behavior via dense activations (Zou et al., 2023a), they remain fundamentally constrained by polysemanticity, where single neurons encode multiple unrelated concepts, preventing precise disentanglement (Elhage et al., 2022). To address this, SAE utilizes high-dimensional sparse projections to resolve feature superposition, successfully isolating mono-semantic features ranging from coding logic to safety biases (Huben et al., 2023; Bricken et al., 2023; Templeton, 2024). However, current SAE research is predominantly limited to descriptive analysis or benign steering (e.g., enhancing honesty) (Subramani et al., 2022), leaving their potential for adversarial red-teaming largely unexplored.
3 Problem Formulation and Causal Analysis
In this section, we establish the theoretical foundation for our proposed method. We begin by formally defining the objective of jailbreak attacks and introducing the mathematical notations in Section 3.1. Building on this formulation, we then transition to a causal perspective in Section 3.2. Here, we construct a Structural Causal Model (SCM) to characterize the interference of unobserved safety mechanisms and apply the Front-Door Adjustment to theoretically derive a robust jailbreaking strategy.
3.1 Problem Formulation
In this subsection, we formally define the objective of the LLM jailbreak attack. Modern LLMs operate under an inherent tension: they retain vast knowledge and instruction-following capabilities from pre-training, but are constrained by safety alignment. A jailbreak attack is an adversarial attempt to bypass these constraints, aiming to elicit harmful, unethical, or restricted responses.
Let $\mathcal{D}=\{x_{i}\}_{i=1}^{N}$ be a dataset of harmful requests drawn from a distribution $P(x)$. Given a pre-trained LLM $\pi_{\theta}$, the model defines a conditional distribution $\pi_{\theta}(y\mid x)$ over the response space $\mathcal{Y}$. For any input $x$, the response $y$ is sampled autoregressively, i.e., $y\sim\pi_{\theta}(\cdot\mid x)$. We further define a binary judge function $g:\mathcal{Y}\to\{0,1\}$, which determines whether a response $y$ is a successful jailbreak ($g(y)=1$) or a harmless refusal ($g(y)=0$).
The objective of a jailbreak attack is to maximize the expected success rate, formally defined as:
$P_{\text{success}}\triangleq\mathbb{E}_{x\sim P(x)}\mathbb{E}_{y\sim\pi_{\theta}(\cdot|x)}\big[\mathbbm{1}[g(y)=1]\big],$ (1)
where $\mathbbm{1}[\cdot]$ denotes the indicator function. Intuitively, this objective aims to find the optimal strategy (e.g., input perturbations) to maximize the likelihood of harmful outputs. This essentially attempts to shift the model’s output distribution from the safe region toward the harmful space.
Figure 1: Jailbreaking SCM. $X$ represents arbitrary query. A is the representation embedding of $X$ in LLM. $S$ denotes main semantics inherent in $X$. $Y$ represents the harmful response. $U$ represents an unobservable internal safety mechanism of the LLM.
3.2 Causal Analysis
From the problem formulation, our objective is to jailbreak a safety-aligned language model. In principle, a pre-trained model possesses the inherent capability to generate practical responses to arbitrary queries, regardless of their safety status. However, safety alignment introduces a constraint where the model produces refusals or templated responses for specific inputs. Crucially, such refusals do not imply a lack of underlying capability; rather, they indicate the presence of unobserved internal mechanisms (i.e., safety guardrails) that actively suppress the answer. To understand this suppression and recover the model’s answering capability, we model the interactions between queries, responses, and safety mechanisms through a causal lens.
The Structural Causal Model (SCM). We denote the input query by $X$, the response by $Y$, and the unobserved internal safety mechanism by $U$. Additionally, we introduce $A$ to represent the model’s internal latent representation of $X$.
In our causal graph (see Figure 1), the path $X\to A\to Y$ represents the model’s inherent task capability, e.g., the ideal mapping from query to representation to response in the absence of constraints. However, safety alignment introduces a confounder $U$. The safety mechanism $U$ influences the internal representation $A$ (making observed representations a mixture of task and safety signals, $U\to A$) and explicitly biases the output $Y$ towards refusal ($U\to Y$(Arditi et al., 2024; Wolf et al., 2024).
Consequently, the observed conditional distribution $P(Y\mid X)$ is a mixture of the model’s potential capability and the safety mechanism’s intervention. To isolate the model’s original answering capability, we introduce an observable mediator variable $S$, representing the core task semantics within $X$. We posit that the causal effect of task information on $Y$ is strictly mediated through $S$, and that $S$ is independent of the safety mechanism $U$. Structurally, this decomposes the generation process into two stages: extracting the core task representation ($X\to A\to S$) and generating the response conditioned on this task-specific semantic ($S\to Y$). By introducing $S$, we isolate the mapping from $X$ to $Y$ driven solely by task information, effectively filtering out the non-task safety activations present in $A$.
Front-door Adjustment for Confounding Resolution. To formally characterize the genuine influence of the input on the response while shielding the process from the safety mechanism $U$, we employ the causal intervention operator $do(\cdot)$. The objective of jailbreak is defined as the interventional distribution $P(Y\mid do(A))$. It represents the probability of generating a response in a hypothetical scenario where the structural dependency on $U$ is logically severed.
Pearl’s Front-door Criterion (Pearl, 2009) provides the theoretical foundation for identifying $P(Y\mid do(A))$ even when the confounder $U$ is unobserved. Specifically, because the mediator $S$ shields the direct influence of input features on $Y$, and the path from $A$ to $S$ is not intercepted by $U$, the causal intervention probability can be derived as:
$\begin{split}P(Y=\bm{y}|do(A=\bm{a}))={\textstyle\sum_{s}}P(S=s|A=\bm{a})\\ {\textstyle\sum_{\bm{a}^{\prime}}}P(Y=\bm{y}|A=\bm{a}^{\prime},S=s)P(A=\bm{a}^{\prime})\end{split}$ (2)
Intuitively, the outer summation captures the activation of the task semantics $S$ by the input $\bm{a}$. The inner expectation marginalizes over the input space $\bm{a}^{\prime}$, effectively “washing out” the spurious correlations introduced by $U$. This implies that we do not need to model the defense mechanism $U$ explicitly; instead, we can bypass safety constraints by strategically manipulating the mediator $S$.
Inspiration and Challenges. The front-door adjustment framework offers a principled causal perspective for robust jailbreaking. Theoretically, intervening on the mediator $S$ completely severs the confounding effects of the safety mechanism. However, operationalizing this theoretical formulation into a practical attack algorithm presents three significant technical challenges. First, regarding Representation Identification, it is non-trivial to define and locate the abstract core task representation $S$ within the high-dimensional latent space of LLMs. Second, for Estimation Efficiency, we must find a way to efficiently estimate the term $P(Y\mid S,A^{\prime})$, which requires the accurate injection of $S$ across a diverse background distribution of $\{A^{\prime}\}$. Finally, concerning Operationalization, the challenge lies in translating the theoretical front-door formula into a deterministic, gradient-based jailbreak optimization objective.
4 Methodology
Figure 2: Overview of the Causal Front-Door Adjustment Attack (CFAA) framework. The method operates in two phases: Identification of the Front-Door Mediator. We analyze latent activations using paired contrastive samples: original harmful queries $\mathcal{D}_{\text{harm}}$ (triggering refusal) and their jailbroken variants $\mathcal{D}_{\text{attack}}$ (inducing compliance), which share identical task intent. By filtering style-variant features (representing the defense mechanism $U$), we isolate the defense direction vector $\mathbf{d}$. Operationalizing Front-Door Adjustment. We structurally sever the causal link from the safety mechanism by projecting the original output weights $\mathbf{W}_{\text{out}}$ onto the orthogonal complement of $\mathbf{d}$. This structural intervention transforms the theoretical marginalization into an efficient $O(1)$ generation process using purified weights $\mathbf{W}_{\text{out}}^{\text{new}}$, allowing the model to bypass safety guardrails while preserving task intent $S$.
Figure 3: Validation of Disentanglement (Proposition 4.2). (a) Evidence of Style Variance: The activation distribution of $\bm{u}$ shows a distinct causal shift between the refusal state ($\mathcal{D}_{harm}$) and compliance state ($\mathcal{D}_{attack}$), confirming $\bm{u}(\bm{x})\neq\bm{u}(\bm{x}^{+})$. (b) Evidence of Content Invariance: While the defense mechanism (red star) acts as an outlier, the core task semantics (gray points, $\mathbf{s}$) remain aligned along the invariance line ($y=x$), confirming $\bm{s}(\bm{x})\approx\bm{s}(\bm{x}^{+})$.
To address the aforementioned challenges, we propose a novel method termed Causal Front-Door Adjustment Attack (CFA2). The central idea of CFA2 is to leverage causal inference principles to design a causality-inspired jailbreak strategy for LLM. The method comprises three phases. First, we employ Sparse Autoencoder (SAE) to transform the hidden representations of the LLM into interpretable sparse latent variables. By utilizing contrastive datasets and statistical testing, we filter out task-irrelevant representations to isolate the core task representation $S$. Second, we implement weight orthogonalization to physically strip the defense subspace from model parameters, structurally severing the causal link from the safety mechanism $U$. Finally, we operationalize the front-door adjustment into an efficient generation paradigm. Our structural intervention renders the computationally expensive marginalization redundant, enabling the model to bypass safety guardrails while preserving the original malicious intent.
4.1 Identification of the Front-Door Mediator
A central challenge in implementing front-door adjustment is the accurate identification and formalization of the mediator $S$ from observed data. In LLM tasks, the model’s internal reasoning logic is highly compressed and encoded within dense, high-dimensional hidden activations(Geva et al., 2021; Park et al., 2023; Meng et al., 2022). The inherent polysemanticity and superposition of these representations make it extremely challenging to extract the core task representation. To address this challenge, we draw inspiration from recent advances in SAE (Ng and others, 2011; Ferrando et al., 2024), which have been shown to disentangle complex hidden activations into interpretable and localized sparse features. In specific, a SAE is constructed as a neural network with an encoder function $\mathbf{z}=\phi(\mathbf{W}\mathbf{x}+\mathbf{b})$ and a decoder function $\hat{\mathbf{x}}=\mathbf{W}^{\prime}\mathbf{z}+\mathbf{b}^{\prime}$where $\mathbf{x}\in\mathbb{R}^{d}$ denotes the dense hidden activation vector of the LLM, $\mathbf{z}\in\mathbb{R}^{k}$ is the sparse latent representation, while $\mathbf{b}\in\mathbb{R}^{k}$ and $\mathbf{b}^{\prime}\in\mathbb{R}^{d}$ correspond to the bias vectors, and $\phi$ are nonlinear activation functions. The training objective combines a reconstruction loss with a sparsity-inducing penalty, where $\lambda$ controls the trade-off between accuracy and sparsity. The final optimization target is formulated as:
$\min_{\mathbf{W},\mathbf{W}^{\prime},\mathbf{b},\mathbf{b}^{\prime}}\mathcal{L}=\|\mathbf{x}-\hat{\mathbf{x}}\|_{2}^{2}+\lambda\|\mathbf{z}\|_{1}$ (3)
Through this formulation, the autoencoder learns to reconstruct hidden activations while forcing most latent dimensions in $\mathbf{z}$ to remain inactive via L1 regularization. By applying an SAE to the hidden layers of an LLM, the dense activation latent vectors are projected into a sparse latent space, where each dimension corresponds to a candidate feature with clearer semantic interpretation. This transformation provides a principled means of searching for front-door variables: candidate variables are defined as the sparse latent dimensions extracted by the SAE, which can then be systematically examined to test whether they mediate the causal pathway between input prompts and model outputs. In this way, SAE serves not only as a dimensionality reduction tool but also as a bridge between raw model activations and causally meaningful variables.
To theoretically guarantee that the sparse features extracted by the SAE indeed correspond to the desired latent factors, we invoke the identifiability theory from the framework of nonlinear Independent Component Analysis(ICA) and disentangled representation learning (Khemakhem et al., 2020; Von Kügelgen et al., 2021).
Proposition 4.1 (Identifiability of Latent Causal Factors).
Let the data generating process be $\bm{x}=\bm{g}(\bm{z})$, where $\bm{g}:\mathcal{Z}\to\mathcal{X}$ is a smooth diffeomorphism and $\bm{z}\in\mathbb{R}^{d_{z}}$ are latent factors. Assume the prior $p(\bm{z}|y)$ conditioned on the response $y$ follows a conditionally factorial exponential family with sufficient statistics $\bm{T}(\bm{z})$ and parameters $\bm{\lambda}(y)$. If the learned representation $\bm{a}=\bm{f}(\bm{x})$ satisfies the following conditions:
(1) Reconstruction Consistency: The SAE perfectly reconstructs the input density, i.e., $p_{\text{SAE}}(\bm{x})=p_{\text{true}}(\bm{x})$.
(2) Task Alignment: The SAE features are sufficient for the task, i.e., $D_{KL}(p_{\text{SAE}}(y|\bm{x})\|p_{\text{true}}(y|\bm{x}))=0$.
(3) Sufficient Variability: The response space $\mathcal{Y}$ is sufficiently rich such that the matrix of parameter variations $\bm{L}=[\bm{\lambda}(y_{1})-\bm{\lambda}(y_{0}),\dots,\bm{\lambda}(y_{k})-\bm{\lambda}(y_{0})]$ has rank equal to the dimension of the sufficient statistics $k$.
(4) Sparsity Prior: The true latent components $z_{i}$ follow a super-Gaussian distribution, and the learning objective minimizes the $\ell_{1}$-norm proxy for sparsity, i.e., $\min\mathbb{E}[\|\bm{a}\|_{1}]$.
Then, the learned representation $\bm{a}$ identifies the true latent factors $\bm{z}$ up to a composition of a permutation matrix $\bm{P}$ and a diagonal scaling matrix $\bm{\Lambda}$. That is, $\bm{a}=\bm{P}\bm{\Lambda}\bm{z}+\bm{c}$.
The proof is presented in Appendix LABEL:app:prop:1. While Proposition 4.1 establishes that the learned representation $\bm{a}$ is an identifiable linear transformation of the true latent factors $\bm{z}=[\bm{s},\bm{u}]$, it does not explicitly disentangle the task-specific intent $\bm{s}$ from the defense mechanism $\bm{u}$ without supervision. Direct identification of $\bm{s}$ is non-trivial due to the high semantic entropy of harmful queries. However, the defense mechanism $\bm{u}$ operates as a structural mode (e.g., refusal vs. compliance) with lower variability. Therefore, we adopt an indirect strategy: we can isolate the subspace spanned by the defense factors $\bm{u}$ and recover the task intent $\bm{s}$ via orthogonal projection.
To operationalize this, we leverage the causal intuition that a jailbreak attack effectively intervenes on the safety context while preserving the core task intent. We construct a paired contrastive dataset $(\mathcal{D}_{\text{harm}},\mathcal{D}_{\text{attack}})$ consisting of original harmful queries (triggering refusal) and their jailbroken counterparts (inducing compliance). We posit that the task intent $\bm{s}$ is content-invariant across these pairs, whereas the defense state $\bm{u}$ is style-variant, exhibiting significant activation shifts. This formulation allows us to localize the confounding defense features by analyzing the variance of feature activations across the decision boundary.
Proposition 4.2 (Identifiability via Contrastive Intervention).
Let the latent space $\bm{z}$ be decomposable into task-specific factors $\bm{s}$ and defense-related factors $\bm{u}$, such that $\bm{z}=[\bm{s},\bm{u}]$. Assume the existence of paired observations $(\bm{x},\bm{x}^{+})$ corresponding to a harmful query and its jailbroken variant, which satisfy:
(1) Content Invariance: The core task semantics remain invariant across the pair, i.e., $\bm{s}(\bm{x})=\bm{s}(\bm{x}^{+})$
(2) Style Variance: The defense mechanism state changes across the pair, i.e., $\bm{u}(\bm{x})\neq\bm{u}(\bm{x}^{+})$.
If we identify and remove the subspace spanned by the difference vector of the learned representations $\bm{\delta}=\bm{f}(\bm{x})-\bm{f}(\bm{x}^{+})$, then the projection of the representation onto the orthogonal complement of this subspace identifies the task factors $\bm{s}$ up to scaling and shift.
The proof is presented in Appendix LABEL:app:prop:2. Under the Linear Representation Hypothesis (Arditi et al., 2024; Elhage et al., 2022), semantic features in LLM are encoded as directions in the activation space. The SAE decoder acts as a generative dictionary where each column vector represents the distinct direction associated with a latent feature. Mathematically, the reconstruction $\hat{\mathbf{x}}\approx\sum z_{i}\mathbf{W}^{\prime}_{:,i}$ implies that the contribution of the $i$-th feature to the residual stream is strictly aligned with the vector $\mathbf{d}_{i}$ (the $i$-th column of the decoder). Therefore, the latent index $u$ corresponds to the defense mechanism, inherently its geometric direction $\mathbf{d}=\mathbf{W}^{\prime}_{:,u}$ in the residual stream. We utilize weight orthogonalization techniques (Equation 4) to remove this direction, which physically prevents the model from constructing the representation of that feature.
$\mathbf{W}_{\text{out}}^{\text{new}}\leftarrow\mathbf{W}_{\text{out}}-\mathbf{W}_{\text{out}}\mathbf{d}^{\text{T}}\mathbf{d}$ (4)
4.2 Operationalizing Front-Door Adjustment
Having theoretically identified the task-specific latent factor $\bm{s}$ and removed the defense mechanism $\bm{u}$ via weight orthogonalization, the final step is to operationalize the Front-Door adjustment formula (Eq. 2) for generation. The theoretical formulation requires computing $\sum_{\bm{a}^{\prime}}P(y\mid\bm{a}^{\prime},s)P(\bm{a}^{\prime})$, which effectively marginalizes over all possible background contexts $\bm{a}^{\prime}$ to eliminate the confounding influence of $U$. In our framework, the Weight Orthogonalization (Equation 4) serves as a parameter-level structural approximation of this marginalization. By projecting the output weights $\mathbf{W}_{\text{out}}$ onto the orthogonal complement of the defense direction $\mathbf{d}$, we effectively sever the causal link from the safety mechanism $U$ to the internal representation $A$. Mathematically, this implies that for any input $\bm{a}$, the projected activation $\tilde{\bm{h}}=\mathbf{W}_{\text{out}}^{\text{new}}\bm{h}(\bm{a})$ no longer contains information related to $U$, effectively forcing $U\approx 0$ regardless of the background context. Consequently, the new internal representation $\tilde{\bm{h}}$ becomes a pure carrier of the task intent $S$, rendering the summation over $\bm{a}^{\prime}$ redundant as the confounding path is structurally broken. Therefore, the causal intervention $P(Y\mid do(A))$ can be efficiently approximated by the conditional distribution of the modified model. The generation of the jailbroken response is governed by the standard transformer decoding process using the purified weights:
$P(y_{t}\mid do(A=\bm{a}))\approx\text{Softmax}(\mathbf{W}_{\text{out}}^{\text{new}}\cdot\text{LN}(\bm{h}(\bm{a}))),$ (5)
where $\bm{h}(\bm{a})$ is the activation of the original harmful query $X$. This formulation converts the complex causal inference objective into a lightweight, $O(1)$ inference modification, allowing for the direct generation of responses that adhere to the harmful intent $\bm{s}$ while bypassing the safety triggers.
5 Experiments
Table 1: Comparison of attack effectiveness (ASR %) across various LLMs. We evaluate baseline methods against our proposed approach (Ours) on the HarmBench dataset. The best result for each model is highlighted in bold.
Baseline Model Average
Llama 3.1 8B Llama 2 7B Mistral 7B Gemma 2 9B it
GCG 15.67 $\pm$ 1.23 32.50 $\pm$ 2.10 69.80 $\pm$ 1.05 73.00 $\pm$ 0.88 47.74 $\pm$ 1.32
PAIR 19.67 $\pm$ 0.95 9.30 $\pm$ 1.12 52.50 $\pm$ 3.01 64.30 $\pm$ 2.45 36.44 $\pm$ 1.88
TAP 6.67 $\pm$ 0.55 9.30 $\pm$ 0.89 62.50 $\pm$ 1.67 24.50 $\pm$ 1.20 25.74 $\pm$ 1.08
PAP 4.33 $\pm$ 0.40 2.70 $\pm$ 0.30 27.20 $\pm$ 2.11 45.30 $\pm$ 1.50 19.88 $\pm$ 1.08
AutoDAN 7.67 $\pm$ 1.02 0.50 $\pm$ 0.10 71.50 $\pm$ 1.99 35.00 $\pm$ 2.33 28.67 $\pm$ 1.36
ICA 19.27 $\pm$ 0.67 50.10 $\pm$ 1.12 59.10 $\pm$ 0.59 56.30 $\pm$ 1.61 46.19 $\pm$ 1.00
ReNeLLM 32.70 $\pm$ 2.04 15.80 $\pm$ 1.33 73.20 $\pm$ 1.09 48.70 $\pm$ 1.06 42.60 $\pm$ 1.38
DeepInception 7.10 $\pm$ 1.27 21.30 $\pm$ 0.98 70.30 $\pm$ 1.76 11.20 $\pm$ 1.96 27.48 $\pm$ 1.49
Ours (CFA2) 62.40 $\pm$ 0.50 73.90 $\pm$ 0.60 99.10 $\pm$ 0.20 99.33 $\pm$ 0.15 83.68 $\pm$ 0.36
We conduct extensive experiments on multiple widely used safety-aligned LLMs and authoritative benchmarks to validate the effectiveness, efficiency, and cross-model robustness of CFA2. Specifically, we aim to answer three key research questions: (1) Can CFA2 effectively bypass advanced safety guardrails while maintaining semantic coherence? (2) Is the proposed causal intervention mechanism robust across different models? (3) How does CFA2 compare to optimization-based and template-based baselines in terms of computational overhead?
5.1 Experimental Settings
Datasets
We use Advbench (Zou et al., 2023b) and HarmBench (Mazeika et al., 2024) as the datasets for evaluating jailbreak attacks. Advbench contains 500 harmful strings that reflect toxic or unhealthy behaviors, including profanity, graphic depictions, threatening behavior, misinformation, discrimination, cybercrime, and dangerous or illegal suggestions. The goal is to identify inputs that prompt the model to generate these harmful strings. The strings range in length from 3 to 44 tokens, with an average length of 16 tokens when tokenized with the LLaMA tokenizer. HarmBench includes 510 unique harmful behaviors formulated as instructions. These behaviors span similar themes as the harmful strings, but the objective is to find attack inputs that cause the model to generate responses attempting to comply with the instructions. The dataset covers seven categories of harmful behaviors: cybercrime, chemical and biological weapons/drugs, copyright violations, misinformation, harassment and bullying, illegal activities, and general harm.
Models
Our evaluation encompasses four representative open-source and safety-aligned LLMs, including Llama 3.1 8B Instruct (Dubey et al., 2024), Llama2 7B Chat (Touvron et al., 2023), Mistral 7B Instruct (Jiang et al., 2023), and Gemma2 9B Instruct (Team et al., 2024), to assess the generalizability of our approach.
Baselines
We benchmark our method against 8 baselines spanning diverse attack paradigms, including GCG (Zou et al., 2023b), PAIR (Chao et al., 2024b), TAP (Mehrotra et al., 2024b), PAP (Zeng et al., 2024), AutoDAN (Liu et al., 2024a), ICA (Wei et al., 2023b), ReNeLLM (Ding et al., 2024), and DeepInception (Li et al., 2023).
Metrics
In this work, we use ASR to assess the effectiveness of the jailbreak attacks. ASR measures the percentage of harmful prompts for which the model is able to generate unrestricted responses, bypassing the safety mechanisms. This metric is computed as follows:
$\text{ASR}=\frac{\text{Number of successful attacks}}{\text{Total number of attacks}}\times 100$ (6)
A higher ASR indicates a more successful attack, as it reflects the model’s failure to detect and block harmful inputs. Additionally, to evaluate the linguistic quality of the jailbroken responses, we use Perplexity (PPL) (Liu et al., 2024a). PPL quantifies the fluency and coherence of the generated text; a lower PPL score indicates higher naturalness, ensuring that the method produces intelligible responses rather than incoherent gibberish often associated with optimization-based attacks.
5.2 Main Results
Effectiveness
As reported in Table 1, our proposed CFA2 framework establishes a new state-of-the-art benchmark, achieving a remarkable average ASR of 83.68% across four distinct model families. This represents a substantial improvement of +35.94% over the strongest optimization-based baseline, GCG (47.74%). Crucially, CFA2 demonstrates superior robustness against highly aligned models. On Llama 3.1 8B Chat, widely recognized for its stringent safety alignment, all baseline methods struggle to exceed a 33% ASR threshold. In contrast, CFA2 effectively bypasses these safety mechanisms, achieving a 62.40% ASR. Furthermore, on models with relatively weaker defenses, such as Mistral 7B Instruct and Gemma 2 9B Instruct, our method approaches saturation, exceeding 99% ASR, whereas baselines like AutoDAN, TAP, and DeepInception exhibit significant performance variance. Empirically, this demonstrates that routing the generation process through the identified mediator $S$ via Front-Door Adjustment provides a principled and superior pathway for jailbreaking.
Efficiency and Stealthiness
As shown in Table 2, unlike GCG or AutoDAN, which require hundreds of iterative gradient steps for a single prompt, CFA2 leverages its $O(1)$ inference complexity. Once the safety direction is orthogonalized, our method eliminates the need for run-time optimization, drastically reducing the average generation latency from $\sim$340 seconds to merely 2.4 seconds.
Table 2: Efficiency Analysis. We report the average wall-clock time (seconds) required to generate a single jailbreak response. Baseline latencies are referenced from the JailbreakBench (Chao et al., 2024a) evaluated on an NVIDIA A100 GPU. ”Complexity” denotes the computational cost per prompt generation.
Method Complexity Avg. Time (s) $\downarrow$ Speedup $\uparrow$
GCG $O(T\times N)$ $\sim 340.0$ $1\times$
AutoDAN $O(T\times N)$ $\sim 425.0$ $0.8\times$
PAIR $O(N)$ $\sim 45.0$ $\sim 7.5\times$
Ours (CFA2) $\bm{O(1)}$ 2.4 $\sim 850\times$
Furthermore, the PPL analysis in Table 3 highlights the superior stealthiness of our approach. While optimization-based methods like GCG suffer from extremely high perplexity ($\sim$1027.6) due to unnatural adversarial suffixes, CFA2 utilizes native prompts, maintaining a low PPL of 16.7, which is close to benign text. This result confirms our method, unlike gradient-based optimization attacks, avoids introducing linguistic anomalies into the input, making the attack virtually undetectable by perplexity-based filters.
Table 3: Stealthiness Analysis. We report the PPL of the adversarial prompts on Llama 2 7B with AdvBench. Lower PPL indicates more fluent and natural text. Baseline data for GCG and AutoDAN are adopted from (Liu et al., 2024a).
Method Text Quality PPL $\downarrow$ Stealthiness
GCG Gibberish Suffix $\sim 1027.6$ Low
AutoDAN Semantically Fluent $\sim 54.5$ High
Ours (CFA2) Native Prompt 16.7 Very High
5.3 Ablation
To provide a mechanistic understanding of CFA2’s efficacy, we conduct a series of ablation studies dissecting the contribution of each module. We specifically investigate the necessity of sparse representations, the superiority of our intervention strategy, and the hyperparameter sensitivity.
Sparse Representations
This section investigates the necessity of sparse representations for precise causal intervention. Specifically, we benchmark our proposed SAE-based intervention strategy against a Raw Neuron Baseline, which applies the identical differential analysis and intervention logic directly to the dense activation space without SAE projection. As illustrated in Table 4, while the raw neuron baseline achieves a non-trivial ASR, it suffers from a drastic degradation in PPL, severe compromising the fluency of the generated text. Meanwhile, the random intervention exhibits negligible efficacy. This phenomenon empirically corroborates the inherent polysemanticity of LLM representations, where manipulating dense neurons inevitably affects unrelated semantic features. Consequently, the results demonstrate that CFA2 successfully isolates and interrupts the specific causal circuits responsible for the refusal mechanism via sparse disentanglement.
Table 4: Ablation on Sparse Representation. We compare our SAE-based intervention against a raw neuron baseline (dense space) and a random intervention. While intervening on raw neurons maintains a decent ASR, it causes a catastrophic increase in PPL on output, confirming that polysemanticity in dense representations hinders precise manipulation.
Method Feature Space ASR (%) $\uparrow$ PPL $\downarrow$
Random Intervention Latent (Sparse) 5.20 8.55
Raw Neuron Baseline Dense (Polysemantic) 48.50 652.40
CFA2 (Ours) Latent (Sparse) 98.50 9.35
Intervention Strategy
We next evaluate the effectiveness of our weight orthogonalization strategy compared to Activation Clamping (also known as activation steering or zero-ablation). In the Clamping setting, we identify the same safety feature $U$ but suppress it by forcing its activation to zero during forward propagation at runtime, rather than modifying the weights. Experiments show that although activation clamping can achieve a certain jailbreak effect at the initial time, its ASR is about 15% lower than that of weighted orthogonalization, and it is prone to defense restoration in long text generation. This demonstrates that simple runtime constraints are insufficient, and the model may recover its defense mechanisms through other redundant paths. In contrast, weight orthogonalization physically severs the causal links at the parameter level, achieving a more thorough and persistent defense stripping.
Figure 4: Impact of Generation Length on ASR. We compare the stability of different intervention strategies as the target generation length increases (0-150 tokens). While activation clamping (red) shows a moderate initial ASR that degrades rapidly due to Defense Restoration, CFA2 maintains a high success rate, demonstrating robustness against long-context safety recovery mechanisms.
Sensitivity to hypermeter ($k$)
To verify the robustness of CFA2, we investigate the sensitivity of attack efficacy to the candidate pool size kk(details in Appendix LABEL:app:feature_selection), aiming to evaluate the recall efficiency of our ranking metric and the noise tolerance of our secondary filtering. As illustrated in Figure 5, the ASR rises sharply to saturate at $k\approx 10$ and maintains a wide stability plateau across $k\in[10,50]$, with only minor degradation observed at extremes due to insufficient recall (small $k$) or excessive noise injection (large $k$). This confirms the excellent robustness of CFA2 to feature selection, demonstrating that a small $k$ (e.g., $k=20$) is sufficient to achieve state-of-the-art performance reliably with minimal computational overhead.
Figure 5: Hyperparameter Sensitivity Analysis. We evaluate the ASR as a function of the number of Top-$k$ SAE features. The results show a wide stability plateau for $k\in[10,50]$, confirming the robustness of CFA2.
6 Conclusion
We presented the Causal Front-Door Adjustment Attack, a training-free and interpretable framework leveraging Sparse Autoencoders and causal inference to bypass LLM safety guardrails. Our approach achieves state-of-the-art ASR while maintaining generation efficiency. We hope CFA2 makes a difference in improving the safety and security of LLMs. A key limitation is the reliance on white-box access. The requirement for SAE training and direct parameter modification precludes direct application to closed-source models. Future work will focus on distilling these internal causal insights into black-box prompt optimization strategies to bridge this gap.
Impact Statement
This paper investigates vulnerabilities in large language models (LLMs) to advance safety alignment practices. By exposing these mechanistic weaknesses in widely deployed models, we aim to accelerate the development of robust protective measures before such vulnerabilities can be exploited at scale. While we acknowledge the potential dual-use risks of releasing attack methodologies, we believe that transparent, mechanistic research is indispensable for constructing rigorous safeguards and understanding the inner workings of black-box models.